6/18 Move PR3
- 6/19: DRMI locked.
- 6/21: polarization issue found.
- 6/21: Stop cooler.
- 6/20: Stop cooler
- 6/28: Open chamber
- Long term:
4/22-5/31: Yarm commissioning
5/27-6/30: DRMI commissioning
- 7/1-8/9: FPMI/DRFPMI commissioning
- 8/12-10/4: ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- Please reduce the number or speed of 16kHz channels.
Report of work
- Type A: ETMY was fixed by opening chamber. Clamp was added to be more tight.
- Needs light to fix position.
- Surface of ITMY was fogged on Sunday. We need to check and wait for a while.
- CRT: closing ETMY after alignment done
- COM: Try to lock PRX for checking beam shape.
- DMG: Try to connect to the additional space on second E18.
- Type A: Oplev of ITMY seems to be something wrong and it will be fixed